OpenStudy (anonymous):

Tickets to a movie cost $9 for each adult and $4.50 for each child. Molly purchased 11 tickets for $72. How many adult and how many child tickets did she purchase? Molly purchased adult tickets and child tickets.

OpenStudy (automaticloveletter):

$9 for each adult $4.50 for each child 11 tickets for $72 Do you have any options?

OpenStudy (anonymous):

I think that this can be solved using simultaneous method, A for Adults and C for Children so 9A + 4.5C = 72 and A + C = 11 A = 11-C substitute 9(11-C) + 4.5C = 72 99 - 9C + 4.5C = 72 -4.5C = -27 C = 6 A = 5 tickets